# HAMLET: Hyperadaptive Agent-based Modeling for Live Embodied Theatrics

### v1

@@ Introduction
Performing Drama
Online performances are based on acts. Each act consists of two components: scene and point. An act can contain one or more scenes and points, which together cover the complete drama process of the act. The scenes define a physical environment in which the drama takes place. It is the stage of ”where” and contains all interactive props. The points define the goals of the plot and are milestones of ”what to do”.
@@
Environment Interaction
Credible physical interaction is the core challenge of the online performance stage. As shown in the Figure 3, we designed a narrator agent to adjudicate all interactions between actors and the environment. The role of the narrator is to ensure the rationality of all physical actions. When an actor attempts to perform a physical action, the narrator will make a judgment based on the environment state and physical rules. If the action is feasible, the narrator will confirm the success, update the environment state, and broadcast an objective description to all participants. Otherwise, the narrator will determine failure with a logical explanation.
Perceive and Decide Module
Hierarchical Control of Performance Effects
All AI actors in the drama performance utilize a layered architecture, which is comprised of an LLM and a PAD module. The LLM is responsible for generating specific speech and actions, while the PAD handles the strategic decision-making that guides them. The design of the PAD is detailed in the following subsection.
@@
In the HAMLET framework, the PAD module plays a crucial role in guiding the AI actors to generate final responses. The design of PAD is conceptually grounded in the dual-process theory of human cognition proposed by Kahneman (Kahneman 2011). According to this theory, human thinking is characterized by two distinct modes of processing: System 1, which is fast, automatic, and intuitive; and System 2, which is slower, deliberate, and analytical. PAD integrates both intuitive and reflective reasoning mechanisms to better simulate human-like decision-making in complex and nuanced drama contexts. Specifically, PAD is mainly responsible for generating a decision of fast, slow, silence or potential actions by tool calls—to simulate and extend the dual-system mechanism. This expanded set of response modes enhances the module’s compatibility with the interactive real-time dramatic context proposed in the study.
Design Principles
As depicted in Figure 4, the core input of PAD is based on a dual principle: subjective and objective perspectives. When perceiving, PAD acts as an AI actor possessing both subjective awareness of its internal state and an objective perception of external stimulus. To ensure well-grounded decision-making, PAD integrates information from both perspectives when a new event occurs.
@@
1:Actor $a_{k}$, Dramatic Context $\mathcal{C}_{drama}$, PAD Model Parameters $\pi_{\theta}$
@@
5:     $\hat{r},[\hat{thinking}]\leftarrow\mathrm{PAD}(prompt_{pad},\pi_{\theta})$
6:     $r\leftarrow\arg\max_{r^{\prime}\in\{\dots\}}P(r^{\prime}|\hat{r})$
@@
Decision-making Process
As detailed in algorithm 1, the decision-making process translates abstract strategic intent into concrete, executable actions through a two-stage process. First, PAD determines a high-level response strategy to set the reaction’s timing and tone. It can also generate the underlying thinking for this choice as an internal monologue through a chain-of-thought process. Subsequently, both the strategy and the generated thinking are used to guide an LLM, which acts as a tool to produce the final, concrete behavior. This behavior is composed of the specific speech to be delivered and a structured action, which is parsed into an executable triplet.
Evaluation Method
Dataset Construction
Our public evaluation dataset was constructed from 100 diverse cases, comprising 50 literary excerpts and 50 custom-authored drama topics. The literary excerpts are selections from 25 classical Chinese [^1] https://m.douban.com/subject_collection/ECKM5FBEI and 25 renowned English [^2] https://sites.prh.com/modern-library-top-100/#top-100-novels literary works. The 50 custom topics, designed to span 10 distinct themes, were developed and reviewed carefully by our annotators. Figure 5 illustrates the thematic distribution of the entire dataset. More details about the public dataset can be found in Appendix A.
@@
Table 1: Performance evaluation of different models based on HAMLETJudge. Bold values indicate the best performance in each column.
Dimension Definition
Defining evaluation dimensions is crucial for quantifying the performance of real-time drama generation. We categorize our evaluation into three core dimensions: Character Performance, Narrative Quality, and Interaction Experience.
Evaluation Principle
Our evaluation principle is to holistically assess dramatic performance, rather than evaluating agent responses individually, turn by turn. This approach is crucial because drama often incorporates literary devices such as foreshadowing and plot twists. Consequently, a seemingly subpar or unusual generation in a single turn might be a deliberate narrative setup for subsequent developments and should not be penalized in isolation.
For the whole drama performance result, we employ HAMLETJudge to automate the evaluation process. This model conducts pairwise comparisons between test result and the baseline result, assigning a score based on a 5-point Likert scale (Robinson 2014) to determine a win rate. The scoring guideline is defined in Appendix B.
Experiments
In this section, we present the experimental results to demonstrate the superiority of our proposed HAMLET framework, alongside ablation studies verifying its reliability and validity.
Settings
To ensure rigorous and reproducible evaluation, we define the baseline and test configurations clearly. All underlying models except the PAD component in HAMLET share the same GPT-4o backbone, with a greedy sampling strategy. Regarding user settings, each agent can be freely configured, and the PAD component is optional.
HAMLET Leaderboard
Following the settings above, we compared various mainstream LLMs ranging from open-source to closed-source and non-reasoning to reasoning models. This reveals their capabilities in English and Chinese online drama performance, serving as a practical reference for real-world applications. The results are shown in Table 1.
Reliability and Validity of HAMLET
Reliability
@@
HAMLETJudge was fine-tuned on pairwise data labeled by annotators. To create this dataset, annotators manually compared paired results and provided judgments with explanations. To validate the effectiveness of our judge model, we measured its agreement with a held-out human validation set using the Pearson correlation coefficient. As shown in Table 2, HAMLETJudge closely aligns with human assessments, outperforming other strong models such as GPT4.1, Claude-4-sonnet and Gemini-2.5-pro.
@@
Table 2: Pearson correlation coefficients (Pearson 1901) of different models and human evaluators.
To further demonstrate the reliability of PAD, we conducted additional experiments evaluating model performance under various response strategies. It is worth noting that, models with stronger general ability tend to have a more homogeneous and consistent performance. Models with output reasoning tokens before tool calls may cause an unbearable time delay in real-time drama. A latency penalty is introduced to capture this effect. The experiment result is shown in Table 3.
Our analysis reveals a distinct performance-latency trade-off among existing models. Models engineered for explicit reasoning achieve balanced performance across strategies but suffer from significant latency penalties. Conversely, non-reasoning models are faster but exhibit a strong performance bias, excelling in Fast scenarios while lacking robustness in complex interactions. In contrast, PAD resolves this problem. It achieves the highest final score by delivering top performance across all strategies, all with zero latency.
Model (FC) Responding Strategy Latency Penalty Final Score
@@
PAD 0.822 0.736 0.711 0 0.756
@@
Validity
@@
In this ablation study, we assess the impact of core design choices by comparing three distinct experimental setups. We randomly select 30 topics from the public evaluation dataset and control the experiment settings as GPT-4o under greedy strategy, then compare: drama performed by (i) raw-prompted GPT-4o only; (ii) the multi-agent collaboration of the full HAMLET framework; and (iii) HAMLET framework with PAD disabled.
Figure 6: Ablation study of HAMLET framework design.
@@
As shown in Figure 6, prompting a single GPT-4o yields substantially lower performance, highlighting the necessity of our multi-agent workflow design. Furthermore, PAD-enabled HAMLET consistently outperforms the version with PAD disabled in all 10 topic categories, demonstrating that PAD serves as a crucial component in the online drama rendition, making the interaction and conversation of AI actors more natural, coherent and humanistic.
Case Study

### v2

@@ 3.2 Online Performance
Performing Drama. Online performances are based on acts. Each act consists of two components: scene and point. An act can contain one or more scenes and points, which together cover the complete drama process. The scenes define a physical environment in which the drama takes place. It is the stage of “where” and contains all interactive props. On the other hand, the points define the goals of the plot and are milestones of “what to do”.
@@
Environment Interaction. Credible physical interaction is the core challenge of the online performance stage. As shown in the Figure 3, we designed a narrator agent to adjudicate all interactions between actors and the environment. The role of the narrator is to ensure the rationality of all physical actions. When an actor attempts to perform a physical action, the narrator will make a judgment based on the environment state and physical rules. If the action is feasible, the narrator will confirm it successful, update the environment state, and broadcast an objective description to all participants. Otherwise, the narrator will determine failure with a logical explanation.
3.3 Perceive and Decide Module
Hierarchical Control of Performance Effects. All AI actors in the drama performance utilize a layered architecture, which is comprised of an LLM and a 0.09804 0.09412 0.23137P0.09804 0.09412 0.23137A0.09804 0.09412 0.23137D\__color_backend_reset:\__color_backend_reset:\__color_backend_reset: module. The LLM is responsible for generating specific speech and actions, while the 0.09804 0.09412 0.23137P0.09804 0.09412 0.23137A0.09804 0.09412 0.23137D\__color_backend_reset:\__color_backend_reset:\__color_backend_reset: handles the strategic decision-making that guides them. The design of the 0.09804 0.09412 0.23137P0.09804 0.09412 0.23137A0.09804 0.09412 0.23137D\__color_backend_reset:\__color_backend_reset:\__color_backend_reset: is detailed in the following subsection.
@@ 3.3 Perceive and Decide Module
In the 0.09804 0.09412 0.23137H0.09804 0.09412 0.23137A0.09804 0.09412 0.23137M0.09804 0.09412 0.23137L0.09804 0.09412 0.23137E0.09804 0.09412 0.23137T\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset: framework, the 0.09804 0.09412 0.23137P0.09804 0.09412 0.23137A0.09804 0.09412 0.23137D\__color_backend_reset:\__color_backend_reset:\__color_backend_reset: module plays a crucial role in guiding the AI actors to generate final responses. The design of 0.09804 0.09412 0.23137P0.09804 0.09412 0.23137A0.09804 0.09412 0.23137D\__color_backend_reset:\__color_backend_reset:\__color_backend_reset: is conceptually grounded in the dual-process theory of human cognition (Kahneman, 2011). According to this theory, human thinking is characterized by two distinct modes of processing: System I, which is fast, automatic, and intuitive; and System II, which is slower, deliberate, and analytical. 0.09804 0.09412 0.23137P0.09804 0.09412 0.23137A0.09804 0.09412 0.23137D\__color_backend_reset:\__color_backend_reset:\__color_backend_reset: integrates both intuitive and reflective reasoning mechanisms to better simulate human-like decision-making in complex and nuanced drama contexts. Specifically, 0.09804 0.09412 0.23137P0.09804 0.09412 0.23137A0.09804 0.09412 0.23137D\__color_backend_reset:\__color_backend_reset:\__color_backend_reset: is mainly responsible for generating a decision of fast, slow, silence or potential actions by tool calls—to simulate and extend the dual-system mechanism. This expanded set of response modes enhances the module’s compatibility with the interactive real-time dramatic context proposed in the study.
Design Principles. As depicted in Figure 4, the input of 0.09804 0.09412 0.23137P0.09804 0.09412 0.23137A0.09804 0.09412 0.23137D\__color_backend_reset:\__color_backend_reset:\__color_backend_reset: is based on a dual principle: subjective and objective perspectives. When perceiving, 0.09804 0.09412 0.23137P0.09804 0.09412 0.23137A0.09804 0.09412 0.23137D\__color_backend_reset:\__color_backend_reset:\__color_backend_reset: acts as an AI actor possessing both subjective awareness of its internal state and an objective perception of external stimulus. To ensure well-grounded decision-making, 0.09804 0.09412 0.23137P0.09804 0.09412 0.23137A0.09804 0.09412 0.23137D\__color_backend_reset:\__color_backend_reset:\__color_backend_reset: integrates information from both perspectives when a new event occurs.
@@
Decision-making Process. As detailed in algorithm 1, the decision-making process translates abstract strategic intent into concrete, executable actions through a two-stage process. First, 0.09804 0.09412 0.23137P0.09804 0.09412 0.23137A0.09804 0.09412 0.23137D\__color_backend_reset:\__color_backend_reset:\__color_backend_reset: determines a high-level response strategy to set the reaction’s timing and tone. It can also generate the underlying thinking for this choice as an internal monologue through a chain-of-thought process. Subsequently, both the strategy and the generated thinking are used to guide an LLM, which acts as a tool to produce the final, concrete behavior. This behavior is composed of the specific speech to be delivered and a structured action, which is parsed into an executable triplet.
@@
1:Actor $a_{k}$, Dramatic Context $\mathcal{C}_{drama}$, 0.09804 0.09412 0.23137P0.09804 0.09412 0.23137A0.09804 0.09412 0.23137D\__color_backend_reset:\__color_backend_reset:\__color_backend_reset: Model Parameters $\pi_{\theta}$
@@
5:  $\hat{r},[\hat{thinking}]\leftarrow\mathrm{{{{0.098040.094120.23137}P{0.098040.094120.23137}A{0.098040.094120.23137}D}\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:}}(prompt_{pad},\pi_{\theta})$
[…]
@@
4 Evaluation Method
Dataset Construction. Our evaluation dataset was constructed from 100 diverse cases, comprising 50 literary excerpts and 50 custom-authored drama topics. The literary excerpts are selections from 25 classical Chinese [^1] https://m.douban.com/subject_collection/ECKM5FBEI and 25 renowned English [^2] https://sites.prh.com/modern-library-top-100/#top-100-novels literary works. The 50 custom topics, designed to span 10 distinct themes, were developed and reviewed carefully by our annotators. Figure 5 illustrates the thematic distribution of the entire dataset. More details about the dataset can be found in Appendix A.
@@ 4 Evaluation Method
Dimension Definition. Defining evaluation dimensions is crucial for quantifying the performance of real-time drama performance. We categorize our evaluation into three core dimensions as follows:
  1. 1.
  2. 2.
  3. 3.
Evaluation Principle. Our evaluation principle is to holistically assess dramatic performance, rather than evaluating agent responses individually, turn by turn. This approach is crucial because drama often incorporates literary devices such as foreshadowing and plot twists. Consequently, a seemingly subpar or unusual generation in a single turn might be a deliberate narrative setup for subsequent developments and should not be penalized in isolation.
For the whole drama performance result, we employ 0.09804 0.09412 0.23137H0.09804 0.09412 0.23137A0.09804 0.09412 0.23137M0.09804 0.09412 0.23137L0.09804 0.09412 0.23137E0.09804 0.09412 0.23137T0.09804 0.09412 0.23137J0.09804 0.09412 0.23137u0.09804 0.09412 0.23137d0.09804 0.09412 0.23137g0.09804 0.09412 0.23137e\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset: to automate the evaluation process. This model conducts pairwise comparisons between test result and the baseline result, assigning a score based on a 5-point Likert scale (Robinson, 2014) to determine a win rate. The detailed scoring guideline is defined in Appendix B.
5 Experiments
In this section, we present the experimental results to demonstrate the superiority of our proposed 0.09804 0.09412 0.23137H0.09804 0.09412 0.23137A0.09804 0.09412 0.23137M0.09804 0.09412 0.23137L0.09804 0.09412 0.23137E0.09804 0.09412 0.23137T\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset: framework, alongside ablation studies verifying its reliability and validity.
Settings. To ensure rigorous and reproducible evaluation, we define the baseline and test configurations clearly. All underlying models except the 0.09804 0.09412 0.23137P0.09804 0.09412 0.23137A0.09804 0.09412 0.23137D\__color_backend_reset:\__color_backend_reset:\__color_backend_reset: component in 0.09804 0.09412 0.23137H0.09804 0.09412 0.23137A0.09804 0.09412 0.23137M0.09804 0.09412 0.23137L0.09804 0.09412 0.23137E0.09804 0.09412 0.23137T\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset: share the same GPT-4o backbone, with a greedy sampling strategy. Regarding user settings, each agent can be freely configured, and the 0.09804 0.09412 0.23137P0.09804 0.09412 0.23137A0.09804 0.09412 0.23137D\__color_backend_reset:\__color_backend_reset:\__color_backend_reset: component is optional.
0.09804 0.09412 0.23137H0.09804 0.09412 0.23137A0.09804 0.09412 0.23137M0.09804 0.09412 0.23137L0.09804 0.09412 0.23137E0.09804 0.09412 0.23137T\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset: Leaderboard. Following the settings above, we compared various mainstream LLMs ranging from open-source to closed-source and non-reasoning to reasoning models. This reveals their capabilities in both English and Chinese online drama performance, serving as a practical reference for real-world applications. The experimental results suggest that reasoning capabilities significantly enhance almost all crucial aspects of the agent’s performance in drama. Notably, specific models such as Claude-4-sonnet-Thinking demonstrated exceptional proficiency in both English and Chinese, establishing them as top contenders for real-world applications that demand high performance in online drama and other role-playing tasks. This model’s balanced and superior performance across all evaluated categories highlights its potential for versatile and effective use in dynamic, interactive environments. The complete result is shown in Table 1.
@@ 5 Experiments
Table 1: Performance evaluation of different models based on 0.09804 0.09412 0.23137H0.09804 0.09412 0.23137A0.09804 0.09412 0.23137M0.09804 0.09412 0.23137L0.09804 0.09412 0.23137E0.09804 0.09412 0.23137T0.09804 0.09412 0.23137J0.09804 0.09412 0.23137u0.09804 0.09412 0.23137d0.09804 0.09412 0.23137g0.09804 0.09412 0.23137e\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:. Bold values indicate the best performance in each column.
5.1 Reliability and Validity of 0.09804 0.09412 0.23137H0.09804 0.09412 0.23137A0.09804 0.09412 0.23137M0.09804 0.09412 0.23137L0.09804 0.09412 0.23137E0.09804 0.09412 0.23137T\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:
@@
Reliability. 0.09804 0.09412 0.23137H0.09804 0.09412 0.23137A0.09804 0.09412 0.23137M0.09804 0.09412 0.23137L0.09804 0.09412 0.23137E0.09804 0.09412 0.23137T0.09804 0.09412 0.23137J0.09804 0.09412 0.23137u0.09804 0.09412 0.23137d0.09804 0.09412 0.23137g0.09804 0.09412 0.23137e\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset: was fine-tuned on pairwise data labeled by annotators. To create this dataset, annotators manually compared paired results and provided judgments with explanations. To validate the effectiveness of our judge model, we measured its agreement with a held-out human validation set using the Pearson correlation coefficient. As shown in Table 2, 0.09804 0.09412 0.23137H0.09804 0.09412 0.23137A0.09804 0.09412 0.23137M0.09804 0.09412 0.23137L0.09804 0.09412 0.23137E0.09804 0.09412 0.23137T0.09804 0.09412 0.23137J0.09804 0.09412 0.23137u0.09804 0.09412 0.23137d0.09804 0.09412 0.23137g0.09804 0.09412 0.23137e\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset: closely aligns with human assessments, outperforming other strong models such as GPT4.1, Claude-4-sonnet and Gemini-2.5-pro.
To further demonstrate the reliability of 0.09804 0.09412 0.23137P0.09804 0.09412 0.23137A0.09804 0.09412 0.23137D\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:, we conducted additional experiments evaluating model performance under various response strategies. It is worth noting that, models with stronger general ability tend to have a more homogeneous and consistent performance. Models with output reasoning tokens before tool calls may cause an unbearable time delay in real-time drama. A latency penalty is introduced to capture this effect. The experiment result is shown in Table 3.
Our analysis reveals a distinct performance-latency trade-off among existing models. Models engineered for explicit reasoning achieve balanced performance across strategies but suffer from significant latency penalties. Conversely, non-reasoning models are faster but exhibit a strong performance bias, excelling in Fast scenarios while lacking robustness in complex interactions. In contrast, 0.09804 0.09412 0.23137P0.09804 0.09412 0.23137A0.09804 0.09412 0.23137D\__color_backend_reset:\__color_backend_reset:\__color_backend_reset: resolves this problem. It achieves the highest final score and stable performance across all strategies with almost zero latency.
@@
Table 2: Pearson correlation coefficients (Pearson, 1901) of different models and human evaluators.
Model (FC)
Responding Strategy
Latency
Penalty
Final
Score
@@
0.09804 0.09412 0.23137P0.09804 0.09412 0.23137A0.09804 0.09412 0.23137D\__color_backend_reset:\__color_backend_reset:\__color_backend_reset: 0.822 0.736 0.711 0 0.756
@@
Figure 6: Ablation study of 0.09804 0.09412 0.23137H0.09804 0.09412 0.23137A0.09804 0.09412 0.23137M0.09804 0.09412 0.23137L0.09804 0.09412 0.23137E0.09804 0.09412 0.23137T\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset: framework design. The radar chart illustrates the comparative evaluation of three experimental configurations in the context of online drama performance.
@@
Validity. In this ablation study, we assess the impact of core design choices by comparing three distinct experimental setups. We randomly select 30 topics from the evaluation dataset and control the experiment settings as GPT-4o under greedy strategy, then compare: drama performed by (i) raw-prompted GPT-4o only; (ii) the multi-agent collaboration of the full 0.09804 0.09412 0.23137H0.09804 0.09412 0.23137A0.09804 0.09412 0.23137M0.09804 0.09412 0.23137L0.09804 0.09412 0.23137E0.09804 0.09412 0.23137T\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset: framework; and (iii) 0.09804 0.09412 0.23137H0.09804 0.09412 0.23137A0.09804 0.09412 0.23137M0.09804 0.09412 0.23137L0.09804 0.09412 0.23137E0.09804 0.09412 0.23137T\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset: framework with 0.09804 0.09412 0.23137P0.09804 0.09412 0.23137A0.09804 0.09412 0.23137D\__color_backend_reset:\__color_backend_reset:\__color_backend_reset: disabled.
@@
As shown in Figure 6, prompting a single GPT-4o yields substantially lower performance, highlighting the necessity of our multi-agent workflow design. Furthermore, 0.09804 0.09412 0.23137P0.09804 0.09412 0.23137A0.09804 0.09412 0.23137D\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:-enabled 0.09804 0.09412 0.23137H0.09804 0.09412 0.23137A0.09804 0.09412 0.23137M0.09804 0.09412 0.23137L0.09804 0.09412 0.23137E0.09804 0.09412 0.23137T\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset: consistently outperforms the version without 0.09804 0.09412 0.23137P0.09804 0.09412 0.23137A0.09804 0.09412 0.23137D\__color_backend_reset:\__color_backend_reset:\__color_backend_reset: in all 10 topic categories, demonstrating that 0.09804 0.09412 0.23137P0.09804 0.09412 0.23137A0.09804 0.09412 0.23137D\__color_backend_reset:\__color_backend_reset:\__color_backend_reset: serves as a crucial component in the online drama rendition, making the interaction and conversation of AI actors more natural, coherent and human-like.
5.2 Case Study